OpenStudy (anonymous):

A florist sells bouquets of daisies and roses. One week, she sold 10 bouquets of roses and 5 bouquets of daisies for $190. The next week, she sold 6 bouquets of roses and 8 bouquets of daisies for $145. Select the equation that best represents this system. Let x represent roses and y represent daisies. 10x + 5y = 145 6x + 8y = 145 5x + 6y = 145 10x + 8y = 190 10x + 6y = 190 5x + 8y = 145 10x + 5y = 190 6x + 8y = 145

OpenStudy (anonymous):

I think you may be looking too into this problem. It's rather simple, as it tells you the equations in the word problem. The first equation is saying that when she sold 10 roses and 5 daisies, she got $190. So, if you put that in math terms it's: 10x+5y=190. The second equation says that she sold 6 roses and 8 daisies for $145. So again in math terms: 6x+8y=145. Hope this helps! (:

OpenStudy (anonymous):

answer: D 10x+5y=190 6x+8y=145